OpenStudy (anonymous):

log(4096) as a fraction

OpenStudy (anonymous):

3.6123599/1

OpenStudy (anonymous):

but as a fraction what would that be

OpenStudy (anonymous):

?=log(64)2

OpenStudy (anonymous):

4^20* (1/128^4)

OpenStudy (anonymous):

I just tried that and it was not correct...it keeps saying it is a fraction

OpenStudy (anonymous):

that is a fraction.... hmmm

OpenStudy (anonymous):

can you explain how to get it...

OpenStudy (accessdenied):

There are plenty of ways to write something as a fraction... Is this under a specific topic on Logarithms like Change of Base?

OpenStudy (anonymous):

It says to evaluate the equation without using a calculator...well I'm kind of cheating because I didn't understand the calculator and I'm on this trying to get help because I'm stuck :(

OpenStudy (anonymous):

like in the example it says log(81)3=12/4...but I don't know how to get the answer to this one

OpenStudy (anonymous):

what is the base of the logarithm?

OpenStudy (anonymous):

64^y=2

OpenStudy (zepp):

\(\huge 64^y=2\) Shoudl be written as \(\large log_{64}2=y\) I think your mistake was to square 64 like this \(\large \log64^2\)

OpenStudy (zepp):

Then get \(\large log(4096)\) Right method to do it is \(\large log_{64}2=y=\frac{1}{6}\)

OpenStudy (zepp):